sara father is fourtimes as old as ahse is. in 18 years he will be exactly twice as old as she. find their present ages
Answer by Cromlix(4381)

Father = 4* Sara's age
F = 4S
In 18 years he will be
exactly twice as old as she is.
F + 18 = 2(S + 18)
Replace F with 4S (from 1st equation)
4S + 18 = 2S + 36
Collect like terms
4S - 2S = 36 - 18
2S = 18
S = 9
Current ages:
Sara = 9 yrs old
Father = 36 years old.
Hope this helps.
:-)